### v2

# GRAFT: GRaPH and Table Reasoning for Textual Alignment — A Benchmark for Structured Instruction Following and Visual Reasoning

Abstract
GRAFT is a structured multimodal benchmark for evaluating models on instruction-following, visual reasoning, and visual-textual alignment tasks. It features programmatically generated charts and synthetically rendered tables, created with Python visualization libraries to ensure control over data semantics, structure, and clarity. Each GRAFT instance pairs a chart or table image with a systematically generated, multi-step analytical question based solely on visual content. Answers are provided in structured formats such as JSON or YAML, supporting consistent evaluation of both reasoning and output format. The benchmark introduces a taxonomy of reasoning types—including comparison, trend identification, ranking, aggregation, proportion estimation, and anomaly detection—to enable comprehensive assessment. Reference answers follow strict factual and formatting guidelines for precise, aspect-based evaluation. GRAFT offers a unified, scalable framework for fine-grained benchmarking of multimodal models on visually grounded, structured reasoning tasks, setting a new evaluation standard in this field.
1 Introduction
Multimodal foundation models are advancing rapidly, excelling in tasks like image captioning, VQA, document parsing, and embodied reasoning. While benchmarks such as VQAv2[2], VizWiz[3], and DocVQA [4] demonstrate progress, they focus mainly on natural images or scanned documents, often lacking structural complexity or controllability. Consequently, these datasets do not fully assess models’ abilities in fine-grained, structured visual reasoning.
Although recent models perform well on general multimodal tasks, their capabilities with structured visual data—such as charts and tables—are not well understood. Tasks involving trends, relationships, or numerical reasoning in such formats require analytical skills and structured outputs (e.g., JSON, YAML), beyond basic perception.
1.1 Motivation
To address this gap, we introduce GRAFT—GRaPH and Table Reasoning for Textual Alignment—a benchmark for evaluating multimodal models on structured reasoning with visual data. GRAFT uses programmatically generated synthetic data, with all visual assets created via Python plotting libraries for precise control and reproducibility.
1.2 Research Questions
GRAFT investigates:
  1. 1.
    How well do current models handle structured visual reasoning across chart and table formats?
  2. 2.
    Can models follow complex instructions and generate correct structured outputs?
  3. 3.
    Do models differ in failure patterns between chart- and table-based inputs?
  4. 4.
    What are the limitations in visual grounding and semantic alignment?
GRAFT thus provides a controlled environment for deeper insights into model reasoning and guides future vision-language research.
The full GRAFT dataset can be found at https://huggingface.co/datasets/ServiceNow-AI/GRAFT_benchmark. We discuss additional dataset details and examples in Appendix C.
2 Related Work
Recent advances in multimodal reasoning span both benchmark development and vision-language model (VLM) architectures. We review (1) the progression of multimodal QA datasets, and (2) instruction-following strategies in VLMs.
2.1 Multimodal Benchmarks
Early benchmarks like VQA v2.0 [5] and COCO-QA [6] addressed basic object and attribute queries, but lacked complex reasoning. Later, synthetic datasets such as ChartQA [7] and PlotQA [8] enabled structured reasoning on charts at scale. Document-based datasets like DocVQA and TAT-DQA [9] focused on text extraction, while GRAFT bridges this with added structural complexity and numerical reasoning, similar to TabFact [10].
2.2 Instruction-Following in Vision-Language Models
Three main strategies have shaped instruction-following in VLMs: (1) task unification (LLaVA [11]), which boosts zero-shot performance through broad instruction-response pairings; (2) visual token compression (Qwen-VL [12]), offering faster inference with minimal accuracy loss; and (3) multi-task alignment (PaLI-X [14]), which reduces forgetting via cross-modal learning. However, instruction-tuned models still lag on multi-hop table reasoning [15]. GRAFT addresses this gap with chained, multi-step reasoning tasks.
3 Methodology
We introduce GRAFT, a fully synthetic benchmark pipeline for evaluating instruction-following and visual reasoning in multimodal large language models (MLLMs) using structured QA over charts and tables. Below, we summarize the data generation and validation process.
Figure 1: Pipeline for image generation and visual reasoning task creation.
3.1 Synthetic Visual Generation
Each GRAFT instance samples domain contexts (e.g., healthcare, finance), assigning persona and scenario metadata.
Charts: Tabular data (8–20 rows, 3–6 columns) is visualized as bar, line, pie, or scatter plots using Matplotlib/Seaborn D.1. LLM-based assessors check type fidelity and label clarity; low-quality charts are filtered out.
Tables: Tabular datasets are synthesized for various formats (pivot, grouped, sortable), then rendered as stylized images using Matplotlib D.3. LLMs assess data fidelity and labels, discarding subpar tables.
3.2 Question and Answer Generation
Approved visuals receive structured descriptions from a VLM, capturing key features. Reasoning-based questions (e.g., ranking, aggregation, anomaly detection) are then generated D.2, D.4. Each QA pair consists of:
  • A question answerable solely from the image.
  • A high-quality answer in JSON or YAML format.
Strict constraints ensure no use of external knowledge, reliance on visual labels for numeric ops, and schema-compliant formatting.
3.3 Jury-of-Judges Filtering
A Jury-of-Judges approach[31] uses multiple LLMs as independent evaluators, each scoring answers by correctness, completeness, visual grounding, and format fidelity.
Consensus strategies included majority as this filtering step ensures high-quality, reliably grounded evaluation data for benchmarking multimodal reasoning in GRAFT.
4 Experimental Setup
We evaluate vision-language models (VLMs) using the GRAFT benchmark, which is divided into two subsets:
  • Chart-QnA: Reasoning over bar, line, scatter, and pie charts.
  • Table-QnA: Analytical tasks with realistic, domain-specific tables.
Each instance provides a rendered chart or table, a natural language question, and a YAML-formatted answer. Tasks include trend analysis, comparison, filtering, aggregation, and multi-step reasoning. Dataset statistics are summarized in Table 1.
Table 1: Summary of GRAFT Benchmark Dataset Statistics
Benchmark Subset #Instances Avg. Q Tokens 90% Q Tokens Avg. A Tokens 90% A Tokens
Chart-QnA 1,412 229 339 164 425
Table-QnA 1,739 323 551 1,022 1,620
Contextual metadata (persona, location, visualization type) is included for each instance, but reasoning must rely solely on the visual input.
4.1 Benchmarking Pipeline
We use a two-stage pipeline (see Figure 2): (1) Prediction Generation: The VLM (e.g., Qwen-2.5-32B-VL) is prompted to generate YAML/JSON answers to each question. (2) Evaluation: Each answer is scored by a GPT4o-based automated judge.
Figure 2: Pipeline for Model Evaluation.
4.2 Evaluation Metrics
Predictions are scored from 1 (poor) to 5 (excellent) along four axes, detailed in Table 2:
  • Correctness: Factual alignment with the visual and reference.
  • Completeness: Fully addresses the query.
  • Visual Grounding: Accurate interpretation of visual elements.
  • Format Fidelity: Strict adherence to structured output format.
Scores are aggregated to summarize model performance, and the judge provides both numerical scores and brief explanations for deviations.
5 Results
We report model performance on GRAFT using our structured, model-agnostic judging pipeline (Section 4.1). Each model output is rated 1–5 on Correctness, Completeness, Visual Grounding, and Format Fidelity; scores are averaged per instance and then across the dataset.
5.1 Quantitative Performance Overview
Figure 3 shows mean scores for Chart-QnA and Table-QnA. Qwen-2.5 32B VL leads both subsets. However, average score alone is insufficient—correctness is most crucial, as errors here undermine the value of other metrics. Our analysis thus prioritizes correctness.
Figure 3: Comparative model performance on Chart-QnA and Table-QnA. Bars represent average scores across all evaluation axes.
For a detailed breakdown of model performance on each subset, see Tables 8 and 9 in the Appendix.
Overall, GRAFT effectively distinguishes between shallow and deep multimodal reasoning abilities. Further discussion is provided in Section F.
6 Conclusion
We introduced the GRAFT Benchmark, a large-scale suite for multimodal question answering over structured tables and charts, designed to test aggregation, comparison, trend analysis, visual grounding, and output fidelity. Using a standardized evaluation pipeline with state-of-the-art vision-language models and LLM-based judging, we found that while models like GPT4o, Pixtral 12B [^1] https://huggingface.co/mistralai/Pixtral-12B-2409, and Qwen 2.5 32B VL perform well on format and completeness, they struggle with correctness and visual grounding—especially for complex graphics. These results highlight the need for improved visual reasoning in multimodal systems. GRAFT provides a robust, extensible platform for evaluating and advancing future multimodal QA models.
7 Limitations and Future Scope
GRAFT has several limitations. First, its reliance on synthetic data may reduce real-world generalizability. Second, language coverage is limited to English, restricting multilingual evaluation. Third, the use of LLM-based judges can sometimes overestimate answer quality, especially in nuanced cases. Finally, some complex chart types—such as heatmaps and sankey diagrams—are underrepresented in the current version.
Future improvements include incorporating human-annotated benchmarks, real-world and multilingual data, richer task formats (e.g., multi-hop reasoning), and finer-grained evaluation of visual grounding. These enhancements will help GRAFT remain a robust and evolving multimodal benchmark.
Appendix A Evaluation Criteria
To ensure consistent and transparent assessment of model responses in GRAFT, we define a scoring matrix across several key evaluation axes. Table 2 outlines the criteria for Correctness, Completeness, Visual Grounding, and Format Fidelity, with detailed interpretations for each score from 1 to 5. This framework facilitates fine-grained, aspect-based evaluation, enabling nuanced analysis of model strengths and weaknesses in structured visual reasoning tasks.
Table 2: Scoring Matrix for Evaluation Metrics (Scale: 1–5)
Evaluation Axis Score Interpretation
Correctness 5 Fully correct and factually aligned with chart/table and reference.
4 One minor factual error with negligible impact.
3 Moderate inaccuracies that weaken reliability.
2 Major factual mistake affecting the core content.
1 Entirely incorrect answer.
Completeness 5 Completely answers all parts of the question.
4 One small aspect is missing.
3 Several required components are missing.
2 Only partially answers the question.
1 Largely incomplete or missing response.
Visual Grounding 5 Perfect interpretation of visual elements.
4 Slight misread but generally correct.
3 Some visual elements misinterpreted.
2 Several major misreads of visual content.
1 No clear connection to the visual data.
Format Fidelity 5 Output matches schema and formatting exactly.
4 Minor format issues (e.g., indentation, punctuation).
3 Multiple format errors, still readable.
2 Major format breakdown (e.g., broken structure).
1 Completely incorrect or unusable format.
Appendix B Benchmark Comparison
To contextualize GRAFT within the broader landscape of multimodal evaluation, Table 3 provides a comparative overview of widely used benchmarks and highlights the unique innovations introduced by GRAFT. For each benchmark, we summarize its core focus and describe how GRAFT advances the field through new capabilities, greater control, or expanded evaluation dimensions. This comparison underscores GRAFT’s role in addressing the limitations of existing datasets and setting a higher standard for structured visual reasoning and output alignment.
Table 3: Comparison of Multimodal Benchmarks and GRAFT’s Innovations
Benchmark Short Description GRAFT’s Improvement
CoSyn [1] Code-guided synthetic multimodal data for text-rich images Provides large-scale synthetic instruction-tuning data; GRAFT complements by targeting fine-grained structured reasoning and strict evaluation protocols
VQA v2.0 [5] Open-ended QA on natural images Adds structured reasoning over charts/tables with schema-constrained outputs
GQA[16] Scene graph QA for compositional reasoning Introduces numerical operations (aggregation, proportion) absent in scene graphs
TextVQA[17] OCR-based QA on text-heavy images Focuses on analytical vs. extraction tasks, with structured answer formats
TallyQA[18] Complex counting with relationships Extends to multi-step reasoning beyond object counting
ChartQA[7] Logical operations on charts Adds programmatic chart generation with pixel-level control
DocVQA[4] Document understanding from scans Shifts focus from text extraction to numerical table/chart analysis
ScienceQA[19] Multimodal science multiple-choice QA Introduces open-ended structured responses (JSON/YAML)
PlotQA[8] 10M synthetic plot QA pairs Adds real-world domain grounding and jury-validated QA pairs
MMBench[20] Comprehensive cyclic evaluation Specializes in fine-grained visual-textual alignment metrics
TAT-DQA[9] Technical document QA Combines textual and visual table reasoning in unified framework
M3SciQA[23] Multi-document scientific QA Focuses on single-document visual reasoning with synthetic data
SceMQA[24] College-level science MCQs Targets analytical vs. recall-based questions with structured outputs
CVR[25] Compositional visual relations Adds real-world domain grounding and format fidelity metrics
Visual Genome[26] Dense image annotations Replaces free-form descriptions with structured analytical tasks
GQA-OOD[27] Out-of-distribution VQA Addresses in-distribution synthetic bias via jury validation
StructMultimodal[28] TextVQA with relation graphs Generalizes to non-OCR structured data (charts/tables)
ChartQA-X[29] Explanations for chart QA Combines answer+explanation generation in structured formats
SPIQA[30] Scientific paper figure QA Focuses on controlled synthetic data vs. real-world scans
Appendix C Dataset Preview
The following appendix provides a visual walkthrough of sample records drawn from the GRAFT benchmark. Each record corresponds to a visual reasoning QA instance—grounded in either a chart or a table—and includes both the visual input and detailed annotations used during training and evaluation.
Each data record is composed of the following key fields:
  • id: A unique SHA-256 hash representing the data instance.
  • image: The rendered chart or table image used as the sole visual input to the model.
  • conversation: A structured list capturing the instruction-response exchange. This includes the visual reference (image) followed by a complex, multimodal question and a YAML-structured answer generated by a VLM.
  • metadata: Encapsulates the quality and validity judgments for each record, including:
    • answerability.JUDGEMENT: Whether the question is visually answerable.
    • answerability.JUDGEMENT _EXPLANATION: Explanation for the answerability tag.
    • correct_answer_evaluation.JUDGEMENT: Whether the answer is factually and visually correct.
    • chosen_formattedness.JUDGEMENT: Assessment of whether the answer adheres to the expected YAML/JSON format.
    • generated_instruction, generated_answer: The actual model-generated instruction and structured output.
    • image_description: Auto-generated natural language description of the visual, capturing axes, legends, trends, and key patterns.
  • judgement_metadata: A consolidated set of scores and explanations from multiple model-based judges (e.g., GPT-4o, Pixtral, Mistral), including their independent verdicts and rationales. This field helps support ensemble-based QA validation pipelines described in Section 4.1.
Each previewed record in the pages below is presented as a compact tabular summary with its associated chart or table image. Fields with long content are truncated for readability (indicated by ellipses).
Sample rendered records are shown below for the first five examples in the benchmark:
Appendix D Instruction Categories
This appendix summarizes the instruction categories used in MIRAGE for generating diverse question types across different visual modalities: graphs, tables, natural images, and posters.
D.1 Graph Generation Taxonomy
Graph-based questions in GRAFT are grounded in high-quality, programmatically rendered charts created from synthetic domain-specific data. The generation process ensures diverse visual styles, realistic semantics, and domain-grounded personas. TableD.3 outlines the key dimensions used for chart synthesis.
Table 4: Taxonomy of Graph Image Generation Attributes
Dimension Description and Examples
Domain Genre Drawn from 25+ application areas including Finance, Education, Environment, and Cybersecurity. Each genre is associated with realistic user roles such as researchers, planners, or policy makers.
Geographic Context Sampled from over 50 global cities and countries to provide cultural and locational grounding. Examples include: Nairobi, Tokyo, San Francisco, and Dubai.
Plot Type Determines visual encoding style. Includes: line, bar, scatter, hist, kde box, heatmap, pie, area, violin count, strip, swarm Each type is selected based on the semantics of the underlying data (e.g., trends, distributions, comparisons).
Visual Styling Rendered using Python plotting libraries (e.g., Matplotlib, Seaborn) with variation in: Axis labels, tick density, color palettes Title fonts and padding DPI scaling (150–350) for clarity and size diversity
Semantic Context Each graph is grounded in a structured metadata schema, which includes: Persona (e.g., high school teacher, financial analyst) Scenario (e.g., monthly earnings report, student performance review) Timeframe and geographic reference
This comprehensive taxonomy ensures the GRAFT graph dataset supports wide-ranging reasoning tasks grounded in diverse, plausible, and richly contextualized scenarios.
D.2 Graph Instruction Categories
Graph-based QA tasks are constructed to probe core visual reasoning capabilities over structured chart imagery. Each question type targets a distinct reasoning skill, grounded solely in the information visible within the chart.
Table 5: Instruction Taxonomy for Graph-Based QA
Category Description Example Question
Comparison Compare values between groups or categories Which group had the highest value?
Trend Identification Identify patterns, slopes, or growth/decline trends What trend is observed from 2010 to 2020?
Change Calculation Determine difference over time or between points By how much did revenue grow from Q1 to Q2?
Ranking Order categories based on visual magnitude Rank regions by number of cases reported
Proportion/Share Determine parts of a whole or segment sizes What share of traffic is from mobile users?
Outlier Detection Identify anomalies or deviations Which year shows an unusual spike in complaints?
Inference Multi-step reasoning from visual data If the current trend continues, when will value exceed 1,000?
Aggregation Sum or combine values from a plot What is the combined total of Category A and B?
Ratio/Rate Compute visual ratios or per-unit rates What is the dropout rate per 100 students?
This instruction taxonomy ensures systematic coverage of reasoning capabilities expected in data visualization tasks and enables benchmarking of visual-language models on structured visual inputs.
D.3 Table Generation Taxonomy
To support structured visual reasoning in the GRAFT benchmark, we design a comprehensive taxonomy for synthetic table generation. This governs not only the content and semantics of tables but also their visual formatting and contextual grounding. Table 6 summarizes the core dimensions of variation used to produce high-fidelity table images.
Table 6: Taxonomy of Table Image Generation Attributes
Dimension Description and Examples
Domain Genre Sampled from 25+ domains (e.g., Healthcare, Finance, Technology) each tied to realistic personas like analysts, doctors, teachers, or planners.
Geographic Context Includes real-world city/country names (e.g., Tokyo, Cape Town, London) to simulate plausible contexts.
Table Type Encodes layout structure. Categories include: flat, grouped, pivot, sortable, highlighted comparison, ranked, time_series, proportion, matrix
Visual Styling Controlled variation in: Font families (e.g., Arial, STIXGeneral), font sizes (8–14pt) Header and row colors, edge borders (black/gray variants) Column/row scaling, zebra striping, padding Render DPI (150–350) for visual fidelity
Layout Controls Additional variation in: Minimum width/height constraints Title font and spacing Highlight rules for emphasizing key cells
This generation taxonomy ensures that each table is semantically grounded, visually distinct, and structurally diverse—reflecting the range of layouts seen in real-world analytical tools and dashboards.
D.4 Table Instruction Categories
Table-based QA tasks are constructed to probe various reasoning capabilities over structured visual data. We define a core taxonomy of instruction categories to ensure analytical diversity across the benchmark. Table7 summarizes these categories with descriptions and example question types.
Table 7: Instruction Taxonomy for Table-Based QA
Category Description Example Question
Comparison Compare values across rows or columns Which group has the highest value?
Pattern Recognition Detect structural or numeric patterns Which column shows a consistent increase?
Change Calculation Compute differences over time or category What changed most between Year X and Y?
Ranking Sort entries based on metrics Rank departments by average score.
Proportion/Share Calculate part-to-whole relationships What % of revenue comes from Product A?
Anomaly Detection Identify unusual highs/lows Which month saw an unexpected drop?
Deductive Inference Multi-step logic or scenario-based reasoning If salaries rise by 10%, what’s the total cost?
Aggregation Sum or average values across a dimension What is the total enrollment in Q3?
Ratio/Rate Derive ratios, per-capita, or unit rates What is the conversion rate per 100 visits?
These QA categories ensure comprehensive coverage of the reasoning spectrum required for structured data interpretation, enabling targeted evaluation of visual understanding and analytical reasoning.
Appendix E Detailed Model Results
E.1 Table-QnA Results
Table 8: Evaluation scores on the Table-QnA subset of GRAFT (1,739 examples).
Model Correctness Completeness Visual Grounding Format Fidelity Avg. Score
Qwen 2.5 32B VL 4.20 4.90 3.85 4.98 4.48
GPT4o 3.78 4.91 3.46 4.97 4.28
Mistral 24B 3.76 4.91 3.44 4.98 4.27
Pixtral 12B 3.40 4.88 3.17 4.99 4.11
GPT4o-mini 3.39 4.80 3.20 4.97 4.09
E.2 Chart-QnA Results
Table 9: Evaluation scores on the Chart-QnA subset of GRAFT (1,412 examples).
Model Correctness Completeness Visual Grounding Format Fidelity Avg. Score
Pixtral 12B 3.92 4.90 3.37 4.99 4.29
Qwen 2.5 32B VL 3.92 4.90 3.34 4.99 4.29
Mistral 24B 3.86 4.94 3.25 4.99 4.26
GPT4o 3.83 4.88 3.25 4.99 4.24
GPT4o-mini 3.62 4.87 3.08 4.99 4.14
Appendix F Model Results Discussion
This section presents a deeper qualitative analysis of model behavior on the GRAFT benchmark, considering architectural, training, and alignment differences. We also integrate insights from our evaluation to highlight where models succeed or fail, and how this reflects their design choices. A synthesis of these findings follows in the form of general takeaways.
F.1 Model-Specific Observations
Qwen-2.5 32B VL demonstrates consistent superiority across all major dimensions—correctness, grounding, and structure. Its high correctness scores across both Chart-QnA and Table-QnA subsets suggest strong internal alignment with step-by-step reasoning, likely supported by extensive instruction tuning and multimodal pretraining. The model’s stable behavior across modalities also indicates robustness in both spatial (charts) and relational (tables) comprehension.
Pixtral 12B is a standout performer on Chart-QnA, attaining the highest visual grounding and average score. However, its performance drops sharply on Table-QnA, particularly in correctness. This dichotomy suggests Pixtral’s strengths lie in shallow visual-text alignment and layout parsing—useful in charts—but it lacks the depth required for multi-step symbolic reasoning over tabular data. This is consistent with its training objectives, which prioritize efficiency and image-level instruction following over structured reasoning.
Mistral-Small-24B-Instruct performs uniformly well across both modalities, especially in format fidelity and completeness. It shows reliable but slightly underwhelming correctness on tables and charts compared to Qwen. Mistral’s performance likely reflects its balanced instruction tuning and decoding stability. However, the absence of a visual modality natively built into Mistral may explain its modest visual grounding scores, suggesting dependence on weak cross-modal adaptation layers.
GPT4o displays strong formatting and completeness scores, which align with OpenAI’s focus on controllability and schema compliance. However, the model consistently underperforms on visual grounding and correctness—especially on complex Chart-QnA questions. This supports the interpretation that GPT4o’s visual capabilities, while promising, are still grounded more in OCR and text layout heuristics than deep visual-semantic parsing. The disparity between visual and textual modalities may account for misalignments in compositional reasoning.
GPT4o-mini, a smaller and more efficient variant, ranks lowest on most axes, particularly correctness and grounding. Despite strong formatting, the model struggles with semantic interpretation of structured data. Its performance illustrates the trade-off between lightweight deployment and high reasoning fidelity. Given its compressed size and likely limited vision training, this result is expected and confirms the importance of scale in visual instruction tuning.
F.2 Analysis and Key Takeaways
These results underscore several crucial observations:
  • Correctness is non-negotiable. No matter how complete or well-formatted a response is, an incorrect answer undermines the purpose of the task. Models with higher correctness scores should be prioritized over those with marginally higher averages but weak reasoning capabilities.
  • Format fidelity is uniformly strong across all models, highlighting the success of schema-constrained instruction templates. This suggests that even smaller models can be guided to produce well-structured outputs when appropriately prompted.
  • Visual grounding remains a bottleneck, particularly for table-based questions that involve implicit relationships, long-range dependencies, or hierarchical reasoning.
  • Instruction tuning and scale matter. Larger instruction-tuned models like Qwen and Mistral outperform others across correctness and grounding, reflecting better task alignment and internal representation.
These insights are further illustrated in the accompanying visual summaries, providing a holistic view of model strengths and limitations across modalities.