### v3

# GRAFT: GRaPH and Table Reasoning for Textual Alignment — A Benchmark for Structured Instruction Following and Visual Reasoning

Abstract
GRAFT is a structured multimodal benchmark for evaluating models on instruction-following, visual reasoning, and visual-textual alignment tasks. It features programmatically generated charts and synthetically rendered tables, created with Python visualization libraries to ensure control over data semantics, structure, and clarity. Each GRAFT instance pairs a chart or table image with a systematically generated, multi-step analytical question based solely on visual content. Answers are provided in structured formats such as JSON or YAML, supporting consistent evaluation of both reasoning and output format. The benchmark introduces a taxonomy of reasoning types—including comparison, trend identification, ranking, aggregation, proportion estimation, and anomaly detection—to enable comprehensive assessment. Reference answers follow strict factual and formatting guidelines for precise, aspect-based evaluation. GRAFT offers a unified, scalable framework for fine-grained benchmarking of multimodal models on visually grounded, structured reasoning tasks, setting a new evaluation standard in this field.
Keywords large language models $\cdot$ multimodal benchmarks $\cdot$ instruction following $\cdot$ visual reasoning $\cdot$ chart understanding $\cdot$ table reasoning $\cdot$ visual-textual alignment $\cdot$ synthetic data generation $\cdot$ evaluation framework
1 Introduction
Recent years have witnessed rapid advancements in multimodal foundation models, which have achieved impressive performance across diverse tasks such as image captioning, visual question answering (VQA), document understanding, and embodied reasoning. Widely used benchmarks, including VQAv2 [2], VizWiz [3], and DocVQA [5], have played a critical role in quantifying progress in these domains. However, these datasets are predominantly oriented toward natural images or scanned documents and often lack either structural complexity or controllable variability. As a result, they provide limited insight into models’ abilities to perform fine-grained reasoning over structured visual content.
Although contemporary multimodal systems excel at broad perceptual and reasoning tasks, their performance on structured visual data—particularly charts and tables—remains insufficiently explored. Reasoning over such inputs requires not only perceptual recognition but also higher-order analytical competencies, such as identifying trends, comparing relationships, and performing numerical inference. Moreover, many practical applications demand structured outputs (e.g., JSON, YAML) that extend beyond natural language responses, thereby introducing further challenges in alignment between visual understanding and structured generation.
1.1 Motivation
To bridge this gap, we present GRAFTGRaPH and Table Reasoning for Textual Alignment—a benchmark explicitly designed to evaluate multimodal models on structured reasoning tasks involving charts and tables. GRAFT leverages programmatically generated synthetic datasets, with all visual assets produced using Python plotting libraries to ensure reproducibility, fine-grained control, and systematic variation in task complexity. This design enables a more rigorous evaluation of reasoning capabilities, disentangled from the noise and biases inherent in natural datasets.
1.2 Research Questions
GRAFT is developed to investigate the following key research questions:
  1. 1.
    To what extent can current multimodal models perform structured reasoning over chart- and table-based inputs?
  2. 2.
    Are models capable of following complex, multi-step instructions and generating semantically and syntactically correct structured outputs?
  3. 3.
    Do error and failure patterns differ systematically between chart reasoning and table reasoning tasks?
  4. 4.
    What are the existing limitations of multimodal models in terms of visual grounding and semantic alignment with textual instructions?
By addressing these questions, GRAFT provides a controlled and transparent environment for diagnosing the reasoning abilities of multimodal systems. The benchmark thereby offers deeper insights into existing limitations and establishes a foundation for advancing vision-language research toward robust structured reasoning.
The full GRAFT dataset is publicly available on Hugging Face [^1] https://huggingface.co/datasets/ServiceNow-AI/GRAFT_benchmark. Additional details regarding dataset construction, along with illustrative examples, are provided in Appendix Dataset Preview.
2 Related Work
Recent advances in multimodal reasoning span both benchmark development and vision-language model (VLM) architectures. We review (1) the progression of multimodal QA datasets, and (2) instruction-following strategies in VLMs.
2.1 Multimodal Benchmarks
Early benchmarks like VQA v2.0 [6] and COCO-QA [7] addressed basic object and attribute queries, but lacked complex reasoning. Later, synthetic datasets such as ChartQA [8] and PlotQA [9] enabled structured reasoning on charts at scale. Document-based datasets like DocVQA and TAT-DQA [10] focused on text extraction, while GRAFT bridges this with added structural complexity and numerical reasoning, similar to TabFact [11].
To contextualize GRAFT within the broader landscape of multimodal evaluation, Table 1 provides a comparative overview of widely used benchmarks and highlights the unique innovations introduced by GRAFT.
Table 1: Comparison of Multimodal Benchmarks and GRAFT’s Innovations
Benchmark Short Description GRAFT’s Improvement
CoSyn [1] Code-guided synthetic multimodal data for text-rich images Provides large-scale synthetic instruction-tuning data; GRAFT complements by targeting fine-grained structured reasoning and strict evaluation protocols
VQA v2.0 [6] Open-ended QA on natural images Adds structured reasoning over charts/tables with schema-constrained outputs
GQA[17] Scene graph QA for compositional reasoning Introduces numerical operations (aggregation, proportion) absent in scene graphs
TextVQA[18] OCR-based QA on text-heavy images Focuses on analytical vs. extraction tasks, with structured answer formats
TallyQA[19] Complex counting with relationships Extends to multi-step reasoning beyond object counting
ChartQA[8] Logical operations on charts Adds programmatic chart generation with pixel-level control
DocVQA[5] Document understanding from scans Shifts focus from text extraction to numerical table/chart analysis
ScienceQA[20] Multimodal science multiple-choice QA Introduces open-ended structured responses (JSON/YAML)
PlotQA[9] 10M synthetic plot QA pairs Adds real-world domain grounding and jury-validated QA pairs
MMBench[21] Comprehensive cyclic evaluation Specializes in fine-grained visual-textual alignment metrics
TAT-DQA[10] Technical document QA Combines textual and visual table reasoning in unified framework
M3SciQA[24] Multi-document scientific QA Focuses on single-document visual reasoning with synthetic data
SceMQA[25] College-level science MCQs Targets analytical vs. recall-based questions with structured outputs
CVR[26] Compositional visual relations Adds real-world domain grounding and format fidelity metrics
Visual Genome[27] Dense image annotations Replaces free-form descriptions with structured analytical tasks
GQA-OOD[28] Out-of-distribution VQA Addresses in-distribution synthetic bias via jury validation
StructMultimodal[29] TextVQA with relation graphs Generalizes to non-OCR structured data (charts/tables)
ChartQA-X[30] Explanations for chart QA Combines answer+explanation generation in structured formats
SPIQA[31] Scientific paper figure QA Focuses on controlled synthetic data vs. real-world scans
2.2 Instruction-Following in Vision-Language Models
Three main strategies have shaped instruction-following in VLMs: (1) task unification (LLaVA [12]), which boosts zero-shot performance through broad instruction-response pairings; (2) visual token compression (Qwen-VL [13]), offering faster inference with minimal accuracy loss; and (3) multi-task alignment (PaLI-X [15]), which reduces forgetting via cross-modal learning. However, instruction-tuned models still lag on multi-hop table reasoning [16]. GRAFT addresses this gap with chained, multi-step reasoning tasks.
3 Methodology
We introduce GRAFT, a fully synthetic benchmark pipeline for assessing instruction-following and structured visual reasoning in multimodal large language models (MLLMs) over charts and tables. The pipeline comprises three stages: (i) synthetic visual generation, (ii) structured QA construction, and (iii) multi-model quality control via a jury-of-judges filter. Figure 1 provides a high-level overview. A complete reference of the variables and notation introduced in this section is provided in Appendix Notation and Variable References.
Figure 1: GRAFT pipeline for synthetic visual generation and structured reasoning task construction.
3.1 Problem Setup and Notation
Let $v$ denote a visual instance (chart or table image), $d(v)$ a structured description of $v$, $q$ a reasoning query grounded in $v$, and $a$ a structured answer adhering to a schema $\mathcal{S}$ (JSON or YAML). An MLLM $f_{\theta}$ receives $(v,q)$ and must produce $\hat{a}\in\mathcal{S}$ such that $\hat{a}$ is semantically correct and visually grounded.
Input Spaces.
Domains $\mathcal{D}$ (e.g., healthcare, finance), personas $\mathcal{P}$, and scenarios $\mathcal{X}$ control content and style; chart types $\mathcal{C}$ and table types $\mathcal{T}$ control structure (Appendices Graph Generation Taxonomy, Table Generation Taxonomy).
Output Targets.
For each instance we produce $(v,d(v),q,a,\mathsf{meta})$, where $\mathsf{meta}$ stores lineage (random seeds, generators, versioned prompts), and validators’ scores.
3.2 Stage I: Synthetic Visual Generation
Inputs.
Domain $d\in\mathcal{D}$, persona $p\in\mathcal{P}$, scenario $x\in\mathcal{X}$; tabular seeds $T\in\mathbb{R}^{m\times n}$ (with $m\in[8,20],\,n\in[3,6]$); chart/table configuration $\kappa$.
Outputs.
A validated visual $v$ (PNG), its rendering specification $\rho$ (Matplotlib/Seaborn parameters), and metadata $\mathsf{meta}_{\text{viz}}$ (axes labels, legends, color maps, font sizes, DPI, seeds).
Procedure.
  1. 1.
    Scenario sampling. Sample $(d,p,x)$ to bind narrative context and label spaces (units, categories).
  2. 2.
    Data synthesis. Generate tabular seed $T$ with typed columns (categorical, ordinal, numerical), injected with controllable traits (trend magnitude, noise level, sparsity) via parameterized generators.
  3. 3.
    Chart rendering. For charts, select $c\in\mathcal{C}$ (bar/line/pie/scatter). Render $v$ via Matplotlib [^2] https://matplotlib.org//Seaborn [^3] https://seaborn.pydata.org/ using $\rho$; enforce axis titles, tick formatting, legends, and numeric precision (Appendix Graph Generation Taxonomy).
  4. 4.
    Table rendering. For tables, select $t\in\mathcal{T}$ (pivot, grouped, sortable); format headers, row bands, and cell alignment; render to image with Matplotlib (Appendix Table Generation Taxonomy).
  5. 5.
    LLM-based visual validation. Use automatic assessors to rate: (i) type fidelity (plot/table matches schema), (ii) label clarity (axes/headers/legend legible, units present), (iii) readability (no occlusion/overplotting). Reject on any failing criterion; otherwise persist $(v,\rho,\mathsf{meta}_{\text{viz}})$.
3.3 Stage II: Structured QA Construction
Inputs.
Validated visual $v$ with spec $\rho$; domain context $(d,p,x)$; instruction-type taxonomies $\Gamma_{\text{chart}}$ and $\Gamma_{\text{table}}$ (e.g., ranking, aggregation, comparison, anomaly detection; Appendices Graph Instruction Categories, Table Instruction Categories); output schema $\mathcal{S}$.
Outputs.
A tuple $(d(v),q,a)$ where $d(v)$ is a structured description, $q$ is a visual-grounded question answerable solely from $v$, and $a\in\mathcal{S}$ is a machine-parseable answer.
Constraints.
(i) No external knowledge: all facts must be derivable from labels or marks in $v$; (ii) Numeric grounding: computations must reference chart/table labels explicitly; (iii) Schema validity: $a$ must satisfy $\mathcal{S}$ (field names, types, enumerations).
Procedure.
  1. 1.
    Saliency extraction. A VLM summarizes $v$ into $d(v)$ capturing encodings (marks, channels), key trends (slopes, extrema), category relations, and plausible aggregations.
  2. 2.
    Instruction sampling. Sample a reasoning template $\gamma\in\Gamma_{\text{chart}}\cup\Gamma_{\text{table}}$ conditioned on $(d,p,x)$ and the structure in $d(v)$.
  3. 3.
    Question synthesis. Instantiate $q$ from $\gamma$ ensuring answerability from $v$ alone; inject clarifying spans (units, time ranges, category subsets) to prevent ambiguity.
  4. 4.
    Answer derivation. Compute $a$ by programmatic reading of $T$ (the generator’s source of truth) and by an independent VLM read from pixels; require agreement. Serialize $a$ to $\mathcal{S}$ (JSON/YAML) with strict formatting.
  5. 5.
    Schema and determinism checks. Validate $a$ against $\mathcal{S}$; re-run derivation with fixed seeds to ensure deterministic reproduction; reject if mismatched.
3.4 Stage III: Jury-of-Judges Filtering
Inputs.
Candidate triples $(v,q,a)$ with $d(v)$ and $\mathsf{meta}$.
Outputs.
A curated set $\mathcal{G}$ of accepted instances; per-instance diagnostics (scores, rationales, failure tags).
Scoring Dimensions.
(i) Correctness (semantic equivalence to ground truth), (ii) Completeness (all requested fields present), (iii) Visual grounding (answer aligns to marks/labels in $v$), (iv) Format fidelity (valid $\mathcal{S}$) [32].
Procedure.
  1. 1.
    Independent evaluation. $K$ LLM judges score each $(v,q,a)$ along the four dimensions with calibrated rubrics (Likert or binary per field) and short textual justifications.
  2. 2.
    Consensus aggregation. Apply majority vote for accept/reject and average for continuous rubrics; tie-break with a deterministic priority (visual grounding $\rightarrow$ correctness $\rightarrow$ format).
  3. 3.
    Adjudication & repair. For borderline cases, attempt auto-repair (format-only or minor rounding) under constraints that forbid semantics changes; re-validate schema; otherwise reject.
  4. 4.
    Finalization. Persist accepted instances to $\mathcal{G}$ with full lineage: $(v,\rho,T,d,p,x,d(v),q,a,\text{scores},\text{judge\_notes},\text{seeds})$.
The GRAFT pipeline combines programmatic data generation, structured reasoning tasks, and rigorous quality control into a unified framework for benchmarking multimodal reasoning. By leveraging synthetic visual assets, schema-constrained QA generation, and multi-model adjudication, GRAFT produces a reproducible and reliable benchmark. Together with the variable reference table in Appendix Notation and Variable References, this methodology provides a transparent foundation for evaluating and diagnosing the structured reasoning capabilities of multimodal models.
4 Experimental Setup
For all experimentation and dataset construction presented in this paper, we employed the SyGra [^4] https://github.com/ServiceNow/SyGra framework [4] as the primary tool for data generation. SyGra’s modular, graph-based architecture enables scalable synthesis, quality tagging, and management of multimodal synthetic data, powering the diverse visual–text datasets used throughout our study.
We evaluate vision-language models (VLMs) using the GRAFT benchmark, which is divided into two subsets:
  • Chart-QnA: Reasoning over bar, line, scatter, and pie charts.
  • Table-QnA: Analytical tasks with realistic, domain-specific tables.
Each instance provides a rendered chart or table, a natural language question, and a YAML-formatted answer. Tasks include trend analysis, comparison, filtering, aggregation, and multi-step reasoning. Dataset statistics are summarized in Table 2.
Table 2: Summary of GRAFT Benchmark Dataset Statistics
Benchmark Subset #Instances Avg. Q Tokens 90% Q Tokens Avg. A Tokens 90% A Tokens
Chart-QnA 1,412 229 339 164 425
Table-QnA 1,739 323 551 1,022 1,620
Contextual metadata (persona, location, visualization type) is included for each instance, but reasoning must rely solely on the visual input.
4.1 Benchmarking Pipeline
We use a two-stage pipeline (see Figure 2): (1) Prediction Generation: The VLM (e.g., Qwen-2.5-32B-VL) is prompted to generate YAML/JSON answers to each question. (2) Evaluation: Each answer is scored by a GPT4o-based automated judge.
Figure 2: Pipeline for Model Evaluation.
4.2 Evaluation Metrics
Predictions are scored from 1 (poor) to 5 (excellent) along four axes, detailed in Table 4:
  • Correctness: Factual alignment with the visual and reference.
  • Completeness: Fully addresses the query.
  • Visual Grounding: Accurate interpretation of visual elements.
  • Format Fidelity: Strict adherence to structured output format.
Scores are aggregated to summarize model performance, and the judge provides both numerical scores and brief explanations for deviations.
5 Discussion
We conduct a qualitative analysis of model performance on GRAFT, focusing on architectural, training, and alignment factors. The goal is to identify where models succeed or fail, and how these outcomes reflect their underlying design choices. We conclude with general takeaways distilled from these observations.
Figure 3: Comparative model performance on Chart-QnA and Table-QnA. Bars represent average scores across all evaluation axes.
5.1 Model-Specific Observations
Qwen-2.5 32B VL[38] achieves the most consistent performance across correctness, grounding, and structure. Its strong results on both Chart-QnA and Table-QnA indicate robust reasoning alignment, likely driven by extensive multimodal pretraining and instruction tuning. The stability across modalities suggests well-integrated spatial (chart) and relational (table) representations.
Pixtral 12B[39] excels on Chart-QnA, obtaining the highest visual grounding scores, but its accuracy degrades on Table-QnA. This divergence suggests that Pixtral’s strengths lie in shallow visual-text alignment and layout parsing—sufficient for charts—but it lacks the symbolic reasoning depth required for tables. The pattern aligns with training objectives optimized for image-level efficiency rather than structured multi-step reasoning.
Mistral-Small-24B-Instruct [^5] https://huggingface.co/mistralai/Mistral-Small-24B-Instruct-2501 delivers balanced performance across modalities, with particular strength in format fidelity and completeness. Its correctness scores, while solid, lag behind Qwen. The absence of a native visual backbone may explain weaker grounding, with performance likely mediated by cross-modal adaptation layers rather than deeply integrated vision modules.
GPT4o[37] demonstrates high format compliance and completeness, reflecting strong alignment with schema-driven prompting. However, it underperforms on correctness and visual grounding, especially for Chart-QnA. This suggests reliance on OCR-style heuristics rather than genuine visual-semantic parsing, which limits compositional reasoning across visual and textual modalities.
GPT4o-mini, the smaller variant, ranks lowest overall, particularly on correctness and grounding. While it maintains schema fidelity, it struggles with semantic interpretation of structured data. These results illustrate the trade-off between deployment efficiency and reasoning capacity, reinforcing the role of scale in effective multimodal alignment.
5.2 Key Takeaways
Our analysis highlights four central findings:
  • Correctness is paramount. Format fidelity and completeness are insufficient if the underlying reasoning is flawed. Models that consistently produce correct answers should be prioritized over those with superficially higher aggregate scores.
  • Schema adherence is largely solved. All models reliably produce structured outputs under schema constraints, indicating that prompt-based guidance is effective even for smaller models.
  • Visual grounding remains the main bottleneck. Table-based reasoning in particular exposes weaknesses in capturing implicit relations, hierarchical structures, and long-range dependencies.
  • Instruction tuning and scale drive performance. Larger, instruction-tuned models (e.g., Qwen, Mistral) achieve better correctness and grounding, underscoring the importance of both capacity and targeted alignment in structured reasoning tasks.
Together, these findings outline the current frontier of multimodal reasoning: while structural compliance is achievable, correctness and grounding—especially over relational data—remain unresolved challenges.
6 Results
We report model performance on GRAFT using our structured, model-agnostic judging pipeline (Section 4.1). Each model output is rated 1–5 on Correctness, Completeness, Visual Grounding, and Format Fidelity; scores are averaged per instance and then across the dataset.
6.1 Quantitative Performance Overview
Figure 3 shows mean scores for Chart-QnA and Table-QnA. Qwen-2.5 32B VL leads both subsets. However, average score alone is insufficient—correctness is most crucial, as errors here undermine the value of other metrics. Our analysis thus prioritizes correctness.
For a detailed breakdown of model performance on each subset, see Tables 9 and 10 in the Appendix.
Overall, GRAFT effectively distinguishes between shallow and deep multimodal reasoning abilities. Further discussion is provided in Section 5.
7 Conclusion
We introduced the GRAFT Benchmark, a large-scale suite for multimodal question answering over structured tables and charts, designed to test aggregation, comparison, trend analysis, visual grounding, and output fidelity. Using a standardized evaluation pipeline with state-of-the-art vision-language models and LLM-based judging, we found that while models like GPT4o, Pixtral 12B [^6] https://huggingface.co/mistralai/Pixtral-12B-2409, and Qwen 2.5 32B VL perform well on format and completeness, they struggle with correctness and visual grounding—especially for complex graphics. These results highlight the need for improved visual reasoning in multimodal systems. GRAFT provides a robust, extensible platform for evaluating and advancing future multimodal QA models.
8 Limitations and Future Scope
GRAFT has several limitations. First, its reliance on synthetic data may reduce real-world generalizability. Second, language coverage is limited to English, restricting multilingual evaluation. Third, the use of LLM-based judges can sometimes overestimate answer quality, especially in nuanced cases. Finally, some complex chart types—such as heatmaps and sankey diagrams—are underrepresented in the current version.
Future improvements include incorporating human-annotated benchmarks, real-world and multilingual data, richer task formats (e.g., multi-hop reasoning), and finer-grained evaluation of visual grounding. These enhancements will help GRAFT remain a robust and evolving multimodal benchmark.